Problem 9-10 Cost of Equity The earnings, dividends, and stock price of Shelby Inc. are expected to grow at 5% per year in thThe earnings, dividends, and stock price of Shelby Inc. are expected to grow at 5% per year in the future. Shelby's common stock sells for $20.50 per share, its last dividend was $1.80, and the company will pay a dividend of $1.89 at the end of the current year.
Answer #1

a) Cost of Equity = Dividend at year end/Price + Growth =1.89/20.50 +5%=14.2195% or 14.22%

b) Cost of equity using CAPM = Risk Free Rate + beta *(Market Return - Risk Free Rate) = 7% +1(13%-7%) = 13.00%

c) Risk Premium in DCF method = 14.2195% - 7% = 7.2195%
Risk Premium in CAPM method = 13% - 7% = 6%
Average risk premium =(7.2195%+6%)/2 = 6.6098%
Cost of equity =Risk Free Rate+Risk Premium =8%+6.6098% =14.6098% or 13.61%

d) Shelby's inc Cost of common equity = (14.2195%+13.00%+14.6098%)/3 = 13.94%